Use the method of compesation to show how you can make the addition of 468+199+399 easier
1
Expert's answer
2019-09-02T05:42:33-0400

468+199+399 = (468 - 1) + (199 + 1) + 399 = 467 + 200 + 399 = (467 - 1) + (399 + 1) + 200 = 466 + 400 + 200 = 466 + 600 = 1066